OpenStudy (saifoo.khan):

Question: 2+3 = 8 3+7 = 27 4+5 = 32 5+8 = 60 6+7 = 72 7+8 = ?

OpenStudy (anonymous):

15

OpenStudy (saifoo.khan):

Wrong^

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

idk

OpenStudy (angela210793):

ur wrong @saifoo.khan :P 2+3=5 :P how did u pass elementary O.o

OpenStudy (anonymous):

^lol

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

so i was right

OpenStudy (saifoo.khan):

WAS*.. but you aren't.

OpenStudy (anonymous):

i want my medal of smartness

OpenStudy (ganpat):

its 98..

OpenStudy (saifoo.khan):

Well, google tells everything.. -_-

OpenStudy (anonymous):

7*6 + 7*8 = ??

OpenStudy (angela210793):

lluminate me...why is 98?

OpenStudy (saifoo.khan):

2+3 = 8 ---------------------------2*(3+1) = 8 3+7 =27.........................3*(7+2) =27 4+5 =32.........................4*(5+3) = 32 5+8 =60 ........................5*(8+4) = 60 6+7 =72........................6*(7+5) =72 7+8 =98 -----------------------7*(8+6) =98 ans.

OpenStudy (ganpat):

the second number is raised by 1 & multiplied by 1st number.. This is for 1st.. similarly raise by 2, .. and so on..

OpenStudy (anonymous):

There is a logic in this: 2*1 + 2*3 = 8 3*2 + 3*7 = 27 4*3 + 4*5 = 32 7*6 + 7*8 = 98..

OpenStudy (angela210793):

tht sounds stupid O.o

OpenStudy (anonymous):

Yes It is..

OpenStudy (saifoo.khan):

lol, really stupid.

OpenStudy (ganpat):

but thats the solution..

OpenStudy (angela210793):

when u say 2+3 why do u have to multiply 2 with (3+1) so stupid -_- :P

OpenStudy (angela210793):

there's no A? O.o